The width of a rectangle is 10 meters and its area is 120 square meters. What is the length of the diagonal?
Answer by nyc_function(2741)

Use A = LW
120 = L(10)
120/10 = L
12 = L
The length is 12 meters.
I will set up the Pythagorean Theorem that you need to find your diagonal.
Keep in mind that the diagonal is the hypotenuse.
Let d = diagonal
(12)^2 + (10)^2 = (d)^2
Can you finish now?